(cosA+secA) holl square +(sinA+cosec) holl square =7+tan square A +cot square

Grade:10

1 Answers

Arun
6 years ago
Dear Saurabh
 
(Cos A + sec A)² + (sinA + cosecA)²
= cos²A + sec²A + 2 + sin²A + cosec²A +2
= 5 + 1 + tan²A + 1 + cot²A        (Sin²A +cos²A = 1)
= 7 + tan²A + cot²A
 
Regards
Arun (askIITians forum expert)